The plastics industry's market demand is x=120-p and market supply is x=p. The amount of externality imposed by producing x units of plastics is 0.5x2. What is the Pigouvian tax? Enter a number (round to two decimal places).
Answer #1

bolution - Given :- Market demand : X = 120-P Market supply :- X=p. oo Marginal poirate Benefit (MPB) = Marginal social Benefusina uthagoras theorema x= -5 IV 4- 4x1*(-240) 281 =-4 176+960 - 2 :-41976 7136621 x cannot be negative value. So y = 13.621